OpenStudy (anonymous):

Evaluate each expression for the given value(s) of the variable(s) DOWN THERE V

OpenStudy (anonymous):

\[(2/3v)^{-3 } v=9\]

OpenStudy (anonymous):

[(27v^3)/(8)]*v=9 27v^4=72 v^4=8/3 v=(8/3)^(1/4)